The text command is nice.

k_{1}, k_{2} = H ± \sqrt{H^{2} − K}  where  \{ \begin{matrix}K & = \frac{eg − f^{2}}{EG − F^{2}} \\ H & = \frac{Eg − 2F f + Ge}{2(EG − F^{2})} \end{matrix}

Interspersing text happens often in math mode.

g = det(g_{ij}) (3–2) (3–3) g^{kl} = (k, l) entry of the inverse matrix of (g_{ij})